# Efficient Reasoning Models: A Survey

### v1

@@ 3.2.2 Pruning or Quantization Retain Reasoning Ability
Recent work (Srivastava et al., 2025) systematically explores the impact of compression techniques like pruning and quantization on the reasoning capabilities of small language models, which shows that while quantization methods (Frantar et al., 2023) have minimal impact on reasoning performance, pruning approaches (Li et al., 2023b) significantly degrade reasoning abilities. Similarly, When Reasoning Meets Compression Zhang et al. (2025b) presents a comprehensive benchmark of compressed LRMs across various reasoning tasks. It also finds that quantized models retain strong reasoning performance and sometimes even surpass the original model, while aggressive pruning causes performance collapse at moderate sparsity. Furthermore, Quantization Hurts Reasoning? (Liu et al., 2025c) systematically evaluates the impact of quantization on reasoning models. It finds that high-bit (e.g., 8-bit) quantization is nearly lossless, while low-bit settings (e.g., 4-bit) significantly degrade performance, especially on complex tasks. Interestingly, the output length of CoT reasoning remains largely unchanged, except under aggressive quantization or when using small models.
3.2.3 Reinforcement Learning Helps Building Small Language Model
@@
SLM-Foresee (Srivastava et al., 2025) conducted a systematic study on the reasoning abilities of diverse small language models, demonstrating that small language models can exhibit strong reasoning potential. Certain models, such as the Qwen2.5 series (Yang et al., 2024a), even achieve performance comparable to or surpassing some LLMs. Open-RS (Dang & Ngo, 2025) enhanced the reasoning capability of small language models using RL with the GRPO algorithm (Guo et al., 2025) and curated a high-quality mathematical reasoning dataset derived from the s1 dataset (Muennighoff et al., 2025) and DeepScaleR dataset (Luo et al., 2025b). They further develop a cosine reward to control response length effectively. Their 1.5B model, trained on 7K samples within 24 hours on 4×A40 GPUs, achieved performance on benchmarks (e.g., AIME2024, MATH-500) that matches or surpasses models like o1-preview (AI., 2024). SimpleRL-Zoo (Zeng et al., 2025) systematically evaluated the generality of ZeroRL (i.e., an RL paradigm that enables LMs to learn long-chain reasoning with only simple rule-based rewards and no additional supervision). The study proposed several key design strategies for successful ZeroRL training: using simple correctness-based rewards, aligning data difficulty with model capacity, and employing stable RL algorithms like GRPO. Remarkably, verification behavior was observed for the first time in small language models outside the Qwen2.5 series [^3] Most existing works focus exclusively on Qwen2.5 models, whose strong instruction following and self-reflection abilities may skew results., further validating the reasoning potential of small language models. Additionally, DeepScaleR [^4] DeepScaleR is a reasoning project for small language models, code and models are available at: https://github.com/agentica-project/deepscaler (Luo et al., 2025b) leverages iterative scaling of GRPO to extend thinking length (i.e., $8\text{K}\rightarrow 16\text{K}\rightarrow 24\text{K}$), significantly improving performance on math reasoning benchmarks. The 1.5B model, DeepScaleR-1.5B-Preview, surpasses O1-Preview and achieves 43.1% Pass@1 on AIME.
@@ 3.3 Let Decoding More Efficient
In the previous sections, we discussed two main directions for improving reasoning efficiency: shortening the reasoning chain and building smaller reasoning models. However, this section covers strategies to accelerate reasoning during inference, focusing on decoding-stage methods that enhance generation speed without compromising reasoning performance. It begins with techniques to reduce computational overhead during TTS (see Section 3.3.1), followed by an overview of other methods for making reasoning faster, with details provided in Section 3.3.2. These methods are summarized in Table 3.
@@
Efficient sampling ST-BoN training-free $C_{3}$ 2.0 $\times$ (L.)
Efficient sampling FastMCTS training-free $C_{4}$uparrow 1.80% Qwen2.5-7B 1.1 - 3.0 $\times$ (T.)
@@
[…]
@@ 3.3.1 Efficiency for Test-Time Scaling Strategy
Efficient Sampling
During the sampling process, the quality of generated reasoning chains often varies, and low-quality outputs lead to substantial redundant computation. A key challenge lies in how to allocate computation more effectively. A natural solution is to terminate low-quality outputs early. Fast Best-of-N (Sun et al., 2024a) proposes speculative rejection, which halts underperforming candidates based on early-stage partial rewards. ST-BoN (Wang et al., 2025b) adopts early consistency checks to identify and retain high-potential candidates while truncating the rest. Early path evaluation can also be applied to reasoning data synthesis. FastMCTS (Li et al., 2025b) leverages MCTS to build reasoning paths while evaluating quality at each step, allowing for dynamic path adjustment. Another line of work explores predicting the future trajectory to reduce redundancy and improve overall quality. Inspired by Model Predictive Control (Qin & Badgwell, 1997), Ma et al. (2024) proposes Predictive-Decoding, which mitigates the myopic nature of token-level generation in CoT by simulating several future reasoning steps (i.e., foresight trajectories) to reweight the token distribution. Similarly, Mendes & Ritter (2025) trains a value model from the language model’s step-by-step generation dynamics to estimate the utility of intermediate reasoning states and decide whether to proceed. $\phi$-Decoding (Xu et al., 2025a) takes a step further by simulating multiple future paths at each step, clustering them to form a representative distribution and sampling the next step from this estimate.
Beyond token-level sampling, recent efforts have focused on structured sampling strategies within multi-path reasoning frameworks such as ToT and SoT. DPTS (Ding et al., 2025) proposes a Dynamic Parallel Tree Search framework that parallelizes reasoning path generation and dynamically manages cache states, enabling flexible path switching without deep exploration. It also incorporates early path evaluation to prioritize promising branches. Similarly, FETCH (Wang et al., 2025a) improves efficiency by merging semantically similar reasoning states to avoid redundant exploration and applying Temporal Difference (TD) learning (Sutton, 1988) with $\lambda$-return to stabilize verifier scores, reducing unnecessary switching. SGD (Jin et al., 2024c) builds a graph over sub-questions to capture their dependencies in a SoT-based framework and introduces difficulty-aware strategies to enable more efficient and higher-quality parallel decoding.
Efficient Self-Consistency
@@ 3.3.2 Other Methods for Making Reasoning Faster
Several other strategies aimed at improving reasoning efficiency also exist, such as introducing decoding stage tricks. One common approach is to decompose the original problem into sub-problems, reducing redundant token generation and skipping uninformative reasoning paths. AoT (Teng et al., 2025) constructs a DAG to model the dependencies among initially decomposed sub-problems. It then solves the overall task by iteratively decomposing and merging sub-problems. At each step, the model only processes a simplified version of the problem, reducing unnecessary token usage, minimizing attention overhead, and avoiding memory issues caused by long contexts. DISC (Light et al., 2025) dynamically partitions the problem into sub-steps and applies reward-based dynamic sampling and early stopping for each step to control compute costs, achieving efficient inference. AR (Liu et al., 2025b) decomposes the reasoning process into atomic reasoning actions organized into an atomic tree and performs structured reasoning via cognitive routing (e.g., reflection, backtracking, and termination). This atomic reasoning paradigm has also proven effective in multimodal large language models (MLLMs) (Xiang et al., 2025).
@@
In real-world applications, we expect LLMs to adjust their output length based on input complexity automatically. For complex problems, the model should allocate more compute to generate a thorough reasoning chain, while for simpler ones, a concise and correct response is preferred. Several methods have been proposed to achieve this. TTC-Optimal Scaling (Snell et al., 2024) proposes a test-time compute-optimal scaling strategy that first estimates the difficulty of a prompt (i.e., either via oracle or model-predicted difficulty) and then adaptively selects different TTS strategies. For instance, on easy questions where the initial response is likely close to correct, self-verification is more efficient than multiple sampling; for complex problems, tree search with a verifier helps explore diverse reasoning paths. MRT (Qu et al., 2025) further improves efficiency by introducing dense rewards based on reasoning progress (i.e., rewarding steps that increase the likelihood of reaching a correct answer) and training LLMs to progress toward solutions and avoid unnecessary computation. RSD (Liao et al., 2025a) enhances efficiency by combining a smaller draft model with a larger target model guided by a reward function. The draft model generates candidate steps, and if the reward is high, the output is accepted; otherwise, the target model refines it. Inspired by meta-cognition (Gao et al., 2024), Meta-Reasoner (Sui et al., 2025b) acts as a strategic advisor to guide the reasoning process, evaluate reasoning progress, and provide high-level guidance (e.g., backtracking, restarting) based on task complexity. Additionally, SpecReason Pan et al. (2025) leverages the semantic tolerance in reasoning processes by using a lightweight model to speculate intermediate steps while reserving the large model for verification and correction.
@@ 4.1.1 General Perspective
  • Memory Footprint refers to the amount of Random Access Memory (RAM) required to run a model during training or inference. This metric is essential for understanding the model’s resource demands, particularly in environments with limited memory capacity, such as edge devices or lightweight servers. Memory is measured in units like MB or GB and is primarily determined by the model size and additional temporary data such as intermediate variables and data structures.
@@ 4.1.2 Reasoning-specific Persective
Additionally, Cuadron et al. (2025) introduced the overthinking score, a reliable metric explicitly designed for quantifying the degree of overthinking in LLMs. The score is obtained using an LLM-based evaluator combined with structured prompt templates. Chen et al. (2024a) proposed the reasoning boundary (RB) to quantify the upper limit of LLM capability in handling complex reasoning tasks (see Appendix A.2.6 for the formula). Wang et al. (2025c) proposed the underthinking metric to evaluate whether a model prematurely abandons effective reasoning paths in incorrect responses, resulting in a large number of unproductive tokens (see Appendix A.2.7 for the formula).
@@ 5 Discussions and Future Directions
While long CoT has been shown to enhance reasoning capabilities, H-CoT (Kuo et al., 2025) reveals that LRMs can be exploited via extended CoT paths to bypass safety guardrails, leading to harmful outputs (Li et al., 2025c). This suggests a tension between safety and efficiency: enhancing safety requires longer, more deliberate reasoning for self-correction, which undermines efficiency, while shorter, efficient reasoning paths may skip critical safety checks. Balancing safety and efficiency thus emerges as a key challenge for future research in LLM reasoning.
@@
Some efficient reasoning methods can be naturally extended to the multimodal large language model (MLLM) setting. The decomposition strategy discussed in Section 3.3.2, which breaks complex tasks into atomic reasoning units, can also benefit multimodal reasoning. Similarly, latent reasoning has shown promise in MLLMs (see Heima in Section 3.1.4). LatentLM (Sun et al., 2024b) further explores this direction by unifying discrete and continuous modalities through latent language modeling. It uses a variational autoencoder (VAE) to encode continuous data into latent vectors and then applies next-token diffusion for autoregressive generation, enabling scalable and efficient multimodal generation.
@@
While long reasoning paths bring remarkable performance, they also cause severe memory issues due to long context. PENCIL (Yang et al., 2025a) addresses this by progressively erasing outdated and unimportant reasoning steps during generation. INFTYTHINK (Yan et al., 2025) adopts a segmentation strategy, breaking the reasoning path into shorter fragments and inserting concise intermediate summaries, enabling chunk-wise thinking. OMNIKV (Hao et al., 2025) observes that adjacent layers share highly similar token importance distributions and thus dynamically select key tokens and reuse them across subsequent layers. MCoT (Yang et al., 2024c) models multi-step reasoning as a Markov chain, where each step depends only on the previous one, avoiding the accumulation of long historical states in the KV cache.
@@
Training long reasoning models remains a computationally intensive task. Recent work has aimed to improve training efficiency through both curriculum learning and RL optimization. Curriculum-based approaches, such as Light-R1 (Wen et al., 2025) and FASTCURL (Song et al., 2025), progressively increase task complexity to facilitate stable learning. Light-R1 employs curriculum SFT and multi-stage post-training, achieving strong performance with public datasets. FASTCURL extends this idea by combining curriculum RL with progressive context window extension, enabling efficient training of R1-like models even on limited hardware. On the RL front, DAPO (Yu et al., 2025b) proposes a scalable and open-source RL system, leveraging decoupled clipping and dynamic sampling for improved training stability. AGPO (Li et al., 2025a) addresses critical instability in the popular GRPO (Guo et al., 2025) by introducing a revised advantage estimation that mitigates zero-variance issues.
Some coreset methods focus on reducing the quantity of training data. LIMO (Ye et al., 2025) argues that complex reasoning abilities are not learned from scratch but elicited through high-quality samples. By constructing a carefully curated dataset of only 817 reasoning samples, the model trained on this data significantly outperforms those trained on nearly 100K examples. The dataset construction involves filtering out easy problems, retaining challenging ones where advanced models struggle, and performing diversity-based sampling. Similarly, s1 (Muennighoff et al., 2025) constructs a compact dataset of 1,000 examples by jointly optimizing for difficulty, diversity, and quality.
@@
Traditional model compression techniques offer valuable opportunities for improving reasoning efficiency. Among them, distillation has demonstrated significant potential in enhancing reasoning efficiency. Distillation effectively transfers reasoning abilities from larger models to smaller ones, enabling them to achieve strong reasoning while significantly reducing costs (see Section 3.2.1). Chen et al. (2025) systematically investigates three key factors that influence the effectiveness of CoT distillation: the granularity of reasoning paths, the format in which reasoning is presented, and the choice of teacher model. These insights offer practical guidance for advancing the distillation of reasoning abilities in small language models. Furthermore, distillation can play a role in other efficient reasoning directions, such as latent reasoning, where it helps compress explicit CoTs into more compact implicit reasoning paths (see Section 3.1.4) and SFT with variable-length CoT data (see Section 3.1.2). Distillation is a promising strategy for efficient reasoning, though there remains room for improvement. Additionally, enhancing the efficiency of the distillation process itself is also a valuable direction for future research. Beyond distillation, other model compression techniques, such as quantization and pruning, also show potential. Although preliminary pruning experiments were not promising, successful quantization suggests that model compression can maintain reasoning performance while improving efficiency in areas like memory usage.

### v2

@@ 3.2.2 Pruning or Quantization Retain Reasoning Ability
Recent work (Srivastava et al., 2025) systematically explores the impact of compression techniques like pruning and quantization on the reasoning capabilities of small language models, which shows that while quantization methods (Frantar et al., 2023b) have minimal impact on reasoning performance, pruning approaches (Li et al., 2023b) significantly degrade reasoning abilities. Similarly, When Reasoning Meets Compression (Zhang et al., 2025b) presents a comprehensive benchmark of compressed LRMs across various reasoning tasks. It also finds that quantized models retain strong reasoning performance and sometimes even surpass the original model, while aggressive pruning causes performance collapse at moderate sparsity. Furthermore, Quantization Hurts Reasoning? (Liu et al., 2025c) systematically evaluates the impact of quantization on reasoning models. It finds that high-bit (e.g., 8-bit) quantization is nearly lossless, while low-bit settings (e.g., 4-bit) significantly degrade performance, especially on complex tasks. Interestingly, the output length of CoT reasoning remains largely unchanged, except under aggressive quantization or when using small models. Notably, the results show that on certain large models, quantization can reduce GPU memory usage by over 75% while retaining nearly 100% of the original performance. Meanwhile, quantized versions of large models are often more effective than standalone small models, offering advantages in both memory efficiency and performance.
3.2.3 Reinforcement Learning Helps Build Small Language Model
@@
SLM-Foresee (Srivastava et al., 2025) conducted a systematic study on the reasoning abilities of diverse small language models, demonstrating that small language models can exhibit strong reasoning potential. Certain models, such as the Qwen2.5 series (Yang et al., 2024a), even achieve performance comparable to or surpassing some LLMs. Open-RS (Dang & Ngo, 2025) enhanced the reasoning capability of small language models using RL with the GRPO algorithm (Guo et al., 2025) and curated a high-quality mathematical reasoning dataset derived from the s1 dataset (Muennighoff et al., 2025) and DeepScaleR dataset (Luo et al., 2025b). They further develop a cosine reward to control response length effectively. Their 1.5B model, trained on 7K samples within 24 hours on 4×A40 GPUs, achieved performance on benchmarks (e.g., AIME 24, MATH-500) that matches or surpasses models like o1-preview (AI., 2024). SimpleRL-Zoo (Zeng et al., 2025a) systematically evaluated the generality of ZeroRL (i.e., an RL paradigm that enables LMs to learn long-chain reasoning with only simple rule-based rewards and no additional supervision). The study proposed several key design strategies for successful ZeroRL training: using simple correctness-based rewards, aligning data difficulty with model capacity, and employing stable RL algorithms like GRPO. Remarkably, verification behavior was observed for the first time in small language models outside the Qwen2.5 series [^2] Most existing works focus exclusively on Qwen2.5 models, whose strong instruction following and self-reflection abilities may skew results., further validating the reasoning potential of small language models. Additionally, DeepScaleR [^3] DeepScaleR is a reasoning project for small language models, code and models are available at: https://github.com/agentica-project/deepscaler (Luo et al., 2025b) leverages iterative scaling of GRPO to extend thinking length (i.e., $8\text{K}\rightarrow 16\text{K}\rightarrow 24\text{K}$), significantly improving performance on math reasoning benchmarks. The 1.5B model, DeepScaleR-1.5B-Preview, surpasses o1-Preview and achieves 43.1% Pass@1 on AIME.
@@ 3.3 Let Decoding More Efficient
In the previous sections, we discussed two main directions for improving reasoning efficiency. However, this section covers strategies to accelerate reasoning during the decoding stage. It begins with techniques to reduce computational overhead during TTS (see Section 3.3.1), followed by an overview of other methods for making reasoning faster, with details provided in Section 3.3.2. These methods are summarized in Table 4, showing that most methods achieve notable efficiency gains and further improve model performance without additional training.
@@
Efficient sampling ST-BoN training-free $C_{3}$ - - 2.0 $\times$ (L.)
Efficient sampling FastMCTS training-free $C_{4}$ $\uparrow$ 1.80% Qwen2.5-7B 1.1 - 3.0 $\times$ (T.)
@@
Other methods AoT training-free - $\uparrow$ 3.00% GPT-4o-mini-0718 -
@@ 3.3.1 Efficiency for Test-Time Scaling Strategy
Efficient Sampling.
During the sampling process, the quality of generated reasoning chains often varies, and low-quality outputs lead to substantial redundant computation. A key challenge lies in how to allocate computation more effectively. A natural solution is to terminate low-quality outputs early. Fast Best-of-N (Sun et al., 2024a) proposes speculative rejection, which halts underperforming candidates based on early-stage partial rewards. ST-BoN (Wang et al., 2025d) adopts early consistency checks to identify and retain high-potential candidates while truncating the rest. Early path evaluation can also be applied to reasoning data synthesis. FastMCTS (Li et al., 2025b) leverages MCTS to build reasoning paths while evaluating quality at each step, allowing for dynamic path adjustment. Another line of work explores predicting the future trajectory to reduce redundancy and improve overall quality. Inspired by Model Predictive Control (Qin & Badgwell, 1997), Ma et al. (2024) proposes Predictive-Decoding, which mitigates the myopic nature of token-level generation in CoT by simulating several future reasoning steps (i.e., foresight trajectories) to reweight the token distribution. Similarly, Mendes & Ritter (2025) trains a value model from the language model’s step-by-step generation dynamics to estimate the utility of intermediate reasoning states and decide whether to proceed. $\phi$-Decoding (Xu et al., 2025a) takes a step further by simulating multiple future paths at each step, clustering them to form a representative distribution and sampling the next step from this estimate.
Beyond token-level sampling, recent efforts have focused on structured sampling strategies within multi-path reasoning frameworks such as ToT and SoT. DPTS (Ding et al., 2025) proposes a Dynamic Parallel Tree Search framework that parallelizes reasoning path generation and dynamically manages cache states, enabling flexible path switching without deep exploration. It also incorporates early path evaluation to prioritize promising branches. Similarly, FETCH (Wang et al., 2025a) improves efficiency by merging semantically similar reasoning states to avoid redundant exploration and applying Temporal Difference (TD) learning (Sutton, 1988) with $\lambda$-return to stabilize verifier scores, reducing unnecessary switching.
Efficient Self-Consistency.
@@ 3.3.2 Other Methods for Making Reasoning Faster
One common approach is to decompose the original problem into sub-problems, reducing redundant token generation and skipping uninformative reasoning paths. AoT (Teng et al., 2025) constructs a DAG to model the dependencies among initially decomposed sub-problems. It then solves the overall task by iteratively decomposing and merging sub-problems. At each step, the model only processes a simplified version of the problem, reducing unnecessary token usage, minimizing attention overhead, and avoiding memory issues caused by long contexts. DISC (Light et al., 2025) dynamically partitions the problem into sub-steps and applies reward-based dynamic sampling and early stopping for each step to control compute costs, achieving efficient inference. AR (Liu et al., 2025b) decomposes the reasoning process into atomic reasoning actions organized into an atomic tree and performs structured reasoning via cognitive routing (e.g., reflection, backtracking, and termination). This atomic reasoning paradigm has also proven effective in multimodal large language models (MLLMs) (Xiang et al., 2025b). SoT (Ning et al., 2023) employs a two-stage decoding strategy by generating a reasoning skeleton and filling nodes in parallel. Inspired by SoT, SGD (Jin et al., 2024c) further builds a graph over sub-questions to capture logical dependencies and introduces difficulty-aware strategies to enable more efficient and higher-quality parallel decoding of reasoning models.
@@
In real-world applications, LLMs are expected to adapt their output length to input complexity, producing detailed reasoning for complex tasks and concise responses for simpler ones. Several methods have been proposed to achieve this. TTC-Optimal Scaling (Snell et al., 2024) proposes a test-time compute-optimal scaling strategy that first estimates the difficulty of a prompt (i.e., either via oracle or model-predicted difficulty) and then adaptively selects different TTS strategies. For instance, on easy questions where the initial response is likely close to correct, self-verification is more efficient than multiple sampling; for complex problems, tree search with a verifier helps explore diverse reasoning paths. MRT (Qu et al., 2025b) further improves efficiency by introducing dense rewards based on reasoning progress (i.e., rewarding steps that increase the likelihood of reaching a correct answer) and training LLMs to progress toward solutions and avoid unnecessary computation. RSD (Liao et al., 2025a) enhances reasoning efficiency by combining a smaller draft model with a larger target model guided by a reward function. The draft model generates candidate steps, and if the reward is high, the output is accepted; otherwise, the target model refines it. Inspired by meta-cognition (Gao et al., 2024), Meta-Reasoner (Sui et al., 2025c) acts as a strategic advisor to guide the reasoning process, evaluate reasoning progress, and provide high-level guidance (e.g., backtracking, restarting) based on task complexity. Additionally, SpecReason (Pan et al., 2025) leverages the semantic tolerance in reasoning processes by using a lightweight model to speculate intermediate steps while reserving the large model for verification and correction.
3.4 A Supplementary: Intersections and Synergies Across Efficient Strategies.
Efficient reasoning strategies are not isolated, many methods combine ideas across categories to achieve better performance and flexibility. Distillation, beyond transferring reasoning capabilities, also serves as an effective means to realize latent reasoning (Deng et al., 2023; Shen et al., 2025c; Yu et al., 2024). Its core idea further supports SFT-based methods by enabling the student model to mimic multi-step reasoning patterns (Kang et al., 2024; Munkhbat et al., 2025). Additionally, SFT and RL can be combined for adaptive reasoning. SFT is used to teach the model different answering modes, while RL helps the model learn when to switch among them based on input difficulty (Fang et al., 2025; Wu et al., 2025b).
@@ 4.1.1 General Perspective
  • Memory Footprint refers to the amount of Random Access Memory (RAM) required to run a model during training or inference. This metric is essential for understanding the model’s resource demands, particularly in environments with limited memory capacity, such as edge devices or lightweight servers. Memory is measured in units like MB or GB and is primarily determined by the model size and additional temporary data (e.g., intermediate variables).
@@ 4.1.2 Reasoning-specific Persective
Additionally, Cuadron et al. (2025) introduced the overthinking score, a reliable metric explicitly designed for quantifying the degree of overthinking in LLMs. The score is obtained using an LLM-based evaluator combined with structured prompt templates. Chen et al. (2024a) proposed the reasoning boundary (RB) to quantify the upper limit of LLM capability in handling complex reasoning tasks (see Appendix A.4.6 for the formula). Wang et al. (2025e) proposed the underthinking metric to evaluate whether a model prematurely abandons effective reasoning paths in incorrect responses, resulting in a large number of unproductive tokens (see Appendix A.4.7 for the formula).
@@ 5 Discussions and Future Directions
While long CoT has been shown to enhance reasoning capabilities, H-CoT (Kuo et al., 2025) reveals that LRMs can be exploited via extended CoT paths to bypass safety guardrails (Feng et al., 2024a), leading to harmful outputs (Li et al., 2025d). This suggests a tension between safety and efficiency: enhancing safety requires longer, more deliberate reasoning for self-correction, which undermines efficiency, while shorter, efficient reasoning paths may skip critical safety checks. Balancing safety and efficiency remains a crucial challenge for future research in LLM reasoning. Latent reasoning offers a more structured, compact, and controllable process, making it a promising direction for reducing safety risks. Additionally, representation alignment, which constrains internal representations, may serve as a lightweight yet effective strategy for enhancing model safety.
@@
Some efficient reasoning methods can be naturally extended to the multimodal large language model (MLLM) setting. The decomposition strategy discussed in Section 3.3.2, which breaks complex tasks into atomic reasoning units, can also benefit multimodal reasoning (Xiang et al., 2025a; Hu et al., 2025). Similarly, latent reasoning has shown promise in MLLMs (see Heima in Section 3.1.4). LatentLM (Sun et al., 2024b) further explores this direction by unifying discrete and continuous modalities through latent language modeling. It uses a variational autoencoder (VAE) to encode continuous data into latent vectors and then applies next-token diffusion for autoregressive generation, enabling scalable and efficient multimodal generation. Additionally, efficient reasoning has been extended to typical vision tasks (Wang et al., 2025c; Koksal & Alatan, 2025; Feng et al., 2025; Li et al., 2025c; Ouyang et al., 2023; Shao et al., 2025), offering valuable insights for future research on integrating structured reasoning into vision-centric multimodal applications.
@@
While long reasoning paths bring remarkable performance, they also cause severe memory issues due to long context. PENCIL (Yang et al., 2025a) addresses this by progressively erasing outdated and unimportant reasoning steps during generation. INFTYTHINK (Yan et al., 2025) adopts a segmentation strategy, breaking the reasoning path into shorter fragments and inserting concise intermediate summaries, enabling chunk-wise thinking. OMNIKV (Hao et al., 2025) observes that adjacent layers share highly similar token importance distributions and thus dynamically select key tokens and reuse them across subsequent layers. MCoT (Yang et al., 2024c) models multi-step reasoning as a Markov chain, where each step depends only on the previous one, avoiding the accumulation of long historical states in the KV cache. These methods show the value of memory-efficient designs; future work should pursue lighter architectures (Gu & Dao, 2024; Yuan et al., 2025) and adaptive context management for scalable long-range reasoning.
@@
Training long reasoning models remains a computationally intensive task. Recent work has aimed to improve training efficiency through both curriculum learning and RL optimization. Curriculum-based approaches, such as Light-R1 (Wen et al., 2025) and FASTCURL (Song et al., 2025), progressively increase task complexity to facilitate stable learning. Light-R1 employs curriculum SFT and multi-stage post-training, achieving strong performance with public datasets. FASTCURL extends this idea by combining curriculum RL with progressive context window extension, enabling efficient training of R1-like models even on limited hardware. On the RL front, DAPO (Yu et al., 2025b) proposes a scalable and open-source RL system, leveraging decoupled clipping and dynamic sampling for improved training stability. AGPO (Li et al., 2025a) addresses critical instability in the popular GRPO (Guo et al., 2025) by introducing a revised advantage estimation that mitigates zero-variance issues. Some coreset methods focus on reducing the quantity of training data. LIMO (Ye et al., 2025) argues that complex reasoning abilities are not learned from scratch but elicited through high-quality samples. By constructing a carefully curated dataset of only 817 reasoning samples, the model trained on this data significantly outperforms those trained on nearly 100K examples. The dataset construction involves filtering out easy problems, retaining challenging ones where advanced models struggle, and performing diversity-based sampling. Similarly, s1 (Muennighoff et al., 2025) constructs a compact dataset of 1,000 examples by jointly optimizing for difficulty, diversity, and quality. Improving training efficiency through algorithmic innovations or data-centric approaches remains a promising direction with substantial room for further exploration.
@@
Traditional model compression techniques offer valuable opportunities for improving reasoning efficiency. Among them, distillation has demonstrated significant potential in enhancing reasoning efficiency. Distillation effectively transfers reasoning abilities from larger models to smaller ones, enabling them to achieve strong reasoning while significantly reducing costs (see Section 3.2.1). Chen et al. (2025b) systematically investigates three key factors that influence the effectiveness of CoT distillation: the granularity of reasoning paths, the format in which reasoning is presented, and the choice of teacher model. These insights offer practical guidance for advancing the distillation of reasoning abilities in small language models. Furthermore, distillation can play a role in other efficient reasoning directions, such as latent reasoning, where it helps compress explicit CoTs into more compact implicit reasoning paths (see Section 3.1.4) and SFT with variable-length CoT data (see Section 3.1.2). Distillation is a promising strategy for efficient reasoning, though there remains room for improvement. Additionally, enhancing the efficiency of the distillation process itself is also a valuable direction for future research. Beyond distillation, other model compression techniques, such as quantization and pruning, also show potential. Although preliminary pruning experiments were not promising, successful quantization suggests that model compression can maintain reasoning performance while improving efficiency in areas like memory usage.
Advancing Sustainability through Efficient Reasoning.
As discussed in this work, efficient reasoning techniques contribute to optimizing the efficiency of reasoning models, reducing computational costs, and minimizing resource usage. These approaches help reduce the carbon footprint by lowering the energy requirements and supporting more environmentally friendly practices. As the use of reasoning models grows, adopting more efficient methods can play a crucial role in mitigating the environmental impact. Additionally, these efficiency improvements do not introduce significant negative effects, ensuring the benefits are realized without unintended consequences.
Comparison with Related Surveys.
Several recent surveys have discussed reasoning models from different angles. For example, Towards Reasoning Era (Chen et al., 2025a) provides a comprehensive overview of long CoT reasoning, focusing primarily on reasoning performance and structure, but does not emphasize efficiency as a central concern. Some surveys (Qu et al., 2025a; Sui et al., 2025b) center on reasoning efficiency. The former (Qu et al., 2025a) organizes methods by stages in the LLM development lifecycle (e.g., pre-training, supervised fine-tuning, reinforcement learning, and inference), offering a broad perspective across the modeling pipeline. The latter (Sui et al., 2025b) classifies approaches based on their core technical mechanisms (e.g., model-based, output-based, and prompt-based), clearly distinguishing the underlying methodological paths. In contrast, our work focuses on how efficiency is achieved during reasoning itself, offering a goal-driven taxonomy centered around making reasoning shorter, smaller, and faster. This structured perspective helps clarify the design space of efficient reasoning and provides clearer guidance for future research.
Connection between Intrinsic Efficiency Metrics and Hard Performance Metrics.
In practical applications, users are primarily concerned with the efficiency that reasoning methods bring to model deployment and usage, typically measured by hard performance metrics such as time and memory. However, efficient reasoning methods often report token count rather than actual runtime. In practice, token count and latency are strongly correlated. We empirically validated this on Qwen2.5-7B using the MAHT-500 dataset, where we observed a clear positive correlation between token count and latency. The Pearson correlation coefficient was 0.9998 with a near-zero p-value, indicating a statistically significant and nearly perfect linear relationship. Meanwhile, some efficient reasoning methods employ PEFT techniques, such as LoRA, to reduce memory usage and calculation costs during the SFT or RL stages. However, this reduction applies only to the training stage and does not affect memory usage during inference or downstream deployment.